- Press Release
- Dec 5, 2022
ISS On-Orbit Status 6 Aug 2002
All ISS systems
continue to function nominally, except as noted previously or
After yesterday’s hard work on EVA preparations, CDR Korzun and
FE-1 Whitson, assisted by FE-2 Treschev, today continued to ready themselves and
their equipment/tools for the two Orlan-suited spacewalks (Russian: VKD) on 8/16
Whitson also had time reserved for a refresher review of DOUG
(dynamic operations ubiquitous graphics) procedures in preparation for
Thursday’s (8/8) MSS (mobile service system) operations. At that time she
will maneuver the SSRMS robotarm to the appropriate position for viewing the
EVA-7 with its Tip LEE, Tip Elbow and Base Elbow video cameras.
[Canadarm2 is currently parked at the POA (payload/ORU accommodation) viewing
position. POA, designed similar to the arm’s LEE (latching end effector),
is a fixture attached to the MBS (mobile base system) for holding future cargoes
as the MBS moves along the truss “railway”. During the last robotics
operations (8/1), POA apparently issued invalid data; the fault event is
currently being analyzed. During Thursday’s robotics ops, Whitson will
first program the arm with auto command settings, so-called Joint OCAS (joint
operator commanded auto sequences), which will then execute the appropriate
joint movements when initiated via laptop. ]
Also in preparation for
her spacewalks, Peggy Whitson completed the obligatory Russian pre-EVA health
assessment test, which essentially uses 30 minutes of vigorous muscle-toning
exercise on the RED (resistive exercise device).
Korzun and Treschev
conducted another two sessions with the Russian-Japanese medical HDTV
(high-definition TV) experiment, that records facial and other crew features
during specific activities on HDTV video for subsequent health
Today was the second day for the current renal (kidney) stone
experiment session, with Valery performing urine collections throughout the day
and Sergei starting out with his dietary/metabolic log entries. Peggy’s
turn begins tomorrow.
It was also time for another one of the periodic
microbiological air samplings on board the station by FE-1 Whitson.
[After powering down the VOA (volatile organic analyzer) for its subsequent
ground-commanded startup and 3.5-hr. calibration run, Whitson deployed the FMK
(formaldehyde monitoring kit) monitors, two in the Lab, two in the Service
Module (SM), for a 24-hour unattended sample collection period. Next, she
set up two SSAS (solid sorbent air sampler) gadgets in Lab and SM for unattended
operation. In sync with the VOA (volatile organic analyzer) performing its
ground-controlled sampling, she also collected an air sample in the Lab with the
GSC (grab sample container), to provide comparison with VOA
Later in the day, Peggy made her weekly entries in the
formal FFQ (food frequency questionnaire) on the MEC (medical equipment
computer) laptop, a requirement of the U.S. HMS (health maintenance systems)
program to monitor U.S. crewmember nutrition.
At 9:35-9:55 am EDT, all
crewmembers participated in a highly enjoyable live interactive educational
event on TV with attendees at the Maryland Science Center. Participants on
the ground included Jim O’Leary (Senior Director of IMAX and Planetarium at the
Maryland Science Center), Roger Crouch (Senior ISS Scientist at NASA HQ),
representatives from Johns Hopkins University & the Space Telescope Science
Institute, middle school students from three Maryland counties, and local
FE-1 Whitson terminated the recharging of the EVA batteries in the
BSA (battery stowage assembly) initiated yesterday. She also started
recharge of the two 110-V batteries used earlier for the TV/Video event. [One
of the EVA battery chargers at first did not detect a voltage for one of the EMU
helmet light batteries when it was inserted into a particular BSA slot.
When swapped with another battery in another slot, the charger read all voltages
on all batteries properly. Battery charging then proceeded
Daily routine maintenance tasks were Sergei Treschev’s
job today (SOSh life support systems inspection, daily Lab payload status check,
and IMS delta file preparation).
At about 10:25 am EDT, the ground began
troubleshooting the still-inoperable MCA (major constituent analyzer) via
S-band, with no crew action required. [The MCA, along with the pressure
control assembly (PCA), trace contaminant control subassembly (TCCS), sample
delivery system (SDS) and carbon dioxide vent valve assembly (CVV), is a
subsystem of the atmosphere control and supply system of the Lab’s environment
control and life support system (ECLSS). Responsibility for the
ECLSS and for its troubleshooting is with the Marshall Space Flight Center
(MSFC) in Huntsville, AL.]
Elektron status: The
Elektron oxygen generator is running. However, the twice-replaced BZh
fluids unit (currently installed: the previous #4 unit) is not functioning
properly due to a malfunctioning upper limit switch for the buffer tank, which
prevents the solenoid valve from closing when the tank is full. After the
crew manually set the valve per ground specialist instructions, the Elektron was
activated and is currently in 32 Amp mode. The valve remains unresponsive
to ground commanding, but Elektron can be powered off without crew intervention
Last Sunday (8/4), the crew completed ground-requested
photo documentation of a closeout panel in the Airlock (A/L), working from the
“job jar” task list. [Expedition 4 crew had noted exposed sharp edges
protruding from a metal screen in the panel’s center. Although the panel
may be replaced with a “beta cloth” closeout containing a ventilation screen as
a permanent solution, the photos will assist engineers in better understanding
of the problem and appropriate interim and permanent
Today’s targets for the CEO (crew earth observation)
program were Industrialized Southeastern Africa (with fair
weather and high pressure over southern Africa, aerosol buildup in the
atmosphere is possible from burning and industrial activity. Of interest on this
pass: looking right of track towards the industrial heartland of South
Africa), Congo-Zimbabwe Biomass Burning (of interest on this
ascending pass: looking left of track over western and northern Zimbabwe for
evidence of fires and smoke plumes and palls in the lower atmosphere, and to try
for oblique and limb views), Lake Poopo (ISS track was directly
over Lake Poopo in good afternoon light; however good context views of the
Altiplano Basin were possible either side of track. Lake Titicaca was to
the north and the bright salars of Uyuni and Coipasa were to the south),
Eastern Mediterranean Dust (a dust storm, visible in satellite
imagery, is moving eastward off the northern coasts of Algeria and
Tunisia. Of interest on this pass as ISS crossed the Libyan coast: looking
left of track towards the Ionian Sea and southern Greece for plumes of dust over
the darker sea surface. On a second pass over the northern Mediterranean,
the crew was to look right of track down the Adriatic Sea for smog palls that
form over Italy at this time of year).
CEO images can be viewed at the
launches (new dates):
Progress 9P (M1-258, or M1-9): 9/20
ISS-9A/STS-112 (S1 truss): NET (not earlier than) 9/28
Soyuz 5S (TMA 211, or TMA-1): 10/28 (approved)
truss): NET 11/2 (under review)
STS-107 (Freestar): NET
U.S. and Russian Segment Status (as of 1:05 pm
Environmental Control and Life Support (ECLSS) and
Thermal Control (TCS):
Elektron O2 generator is powered On (32-amp
mode). Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5 (vacuum
pump failed). U.S. CDRA CO2 scrubber is Off (leak). BMP Harmful
Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify
SM Working Compartment: Pressure (mmHg) — 760,
temperature (deg C) — 26.9, ppO2 (mmHg) — 158.8, ppCO2 (mmHg) — 2.6.
Transfer Compartment: Pressure (mmHg) — 761, temperature (deg C) —
FGB Cabin: Pressure (mmHg) — 756, temperature (deg C) —
Node: Pressure (mmHg) — 753.43, temperature (deg C) —
25.1 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
Lab: Pressure (mmHg) — 755.02, temperature (deg C) — 24.8, ppO2
(mmHg) — n/a; ppCO2 (mmHg) — n/a;
Joint Airlock (Equip. Lock):
Pressure (mmHg) — 755.02, temperature (deg C) — 28.0; shell heater temp (deg
C) — 27.9, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
Shell heater temp (deg C) — 24.2
PMA-2: Shell heater temp (deg
C) — 14.3
(Note: Partial pressures ppO2 and ppCO2 in U.S. segment
[USOS] not available because MCA [major constituent analyzer] is failed and in
Extended Life mode [= a state that preserves mass spectrometer vacuum but
produces no pp data]). MSA (mass spectrometer assembly) and VGA (verification
gas assembly) were replaced, but some more work needs to be
Electrical Power Systems (EPS):
Beta Gimbal Assembly
(BGA) 2B in Autotrack mode; BGA 4B also in Autotrack mode.
Battery #1 is off line, battery #5 is cycling; all other batteries (6) are in
“Partial charge” mode.
FGB batteries: no data available
Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby
Thermal Control Systems:
Air conditioner SKV-1 is
On. SKV-2 is Off.
Command & Data Handling
C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1
is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
LA-1, LA-2 and LA-3
MDMs are all operating.
PL-1 MDM is operational; PL-2 MDM is Off.
(automated payload switch #1) and APS-2 are both On.
SM Terminal Computer
(TVM): 3 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 3
redundant lanes (of 3) operational.
State vector — US GPS (SIGI string 1)
Attitude — Russian
Angular rates — US RGA1 (rate gyro assembly
All Russian communications
& tracking systems are nominal.
S-band is operating nominally.
is operating nominally.
Audio subsystem operating nominally.
subsystem operating nominally.
MCOR (medium-rate communications outage
recorder) is operating
SSRMS/Canadarm2 at MBS PDGF 1
(mobile base system/power & data grapple fixture 1) and MBS PDGF 4, with
Keep Alive power on both strings (based on MBS).
MBS: Operational on
RWS (robotics workstations): Cupola is Active, Lab RWS is
ISS Orbit (as of this morning, 7:39 am EDT [=
Mean altitude — 395.6 km
Apogee — 407.9 km
Period — 92.5 min.
Inclination (to Equator) — 51.64
Eccentricity — 0.0018108
Orbits per 24-hr. day — 15.57
decrease — 75 m (mean) in last 24 hours
Solar Beta Angle — 29.0 deg
Revolutions since FGB/Zarya launch (Nov. ’98)
Current Flight Attitude — XPOP (x-axis perpendicular to orbit
plane = “sun-fixed” [yaw: ~180 deg, pitch: -5.7 deg., roll: 0 deg]).
Next maneuver to LVLH (“earth-oriented”) on Friday, 8/9.
For more on ISS
orbit and worldwide naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html